### v1

# Skyrise: Exploiting Serverless Cloud Infrastructure for Elastic Data Processing

Abstract
Serverless computing offers elasticity unmatched by conventional server-based cloud infrastructure. Although modern data processing systems embrace serverless storage, such as Amazon S3, they continue to manage their compute resources as servers. This is challenging for unpredictable workloads, leaving clusters often underutilized. Recent research shows the potential of serverless compute resources, such as cloud functions, for elastic data processing, but also sees limitations in performance robustness and cost efficiency for long running workloads. These challenges require holistic approaches across the system stack. However, to the best of our knowledge, there is no end-to-end data processing system built entirely on serverless infrastructure. In this paper, we present Skyrise, our effort towards building the first fully serverless SQL query processor. Skyrise exploits the elasticity of its underlying infrastructure, while alleviating the inherent limitations with a number of adaptive and cost-aware techniques. We show that both Skyrise’s performance and cost are competitive to other cloud data systems for terabyte-scale queries of the analytical TPC-H benchmark.
1 Introduction
Organizations increasingly move their data processing workloads to the cloud. One main driver is elasticity, which allows them to deploy virtual resources on demand and avoid the upfront cost of on-premise deployment. In particular, data warehousing and analytics with high and fluctuating resource demands benefit from flexible provisioning. The revenue of cloud-based systems, such as Redshift [1], BigQuery [2], and Snowflake [3] lately surpassed the revenue of on-premise systems [4] and continues to grow rapidly [5].
In the last decade, public cloud infrastructure has evolved to become more elastic. This evolution includes scalable datacenter networking [6, 7], accelerated server virtualization [8], and finer-grained economical models for compute resources [9, 10]. The benefits of these features, however, do not translate to cloud-hosted systems automatically and require architectural changes.
Systems are being rearchitected to exploit the elasticity on the cloud infrastructure level [11, 12, 13]. There are three major design changes: First, cloud systems disaggregate their compute and storage resources to scale them independently. Second, they split their work into small and idempotent tasks that can be scheduled quickly, as well as redundantly in the case of transient errors in the infrastructure. Third, they optimize cost given the economics of their infrastructural building blocks.
Serverless cloud computing is a continuation of this evolution, with most of the related challenges intensifying. Serverless compute units are tiny and stateless [14, 15]. Large-scale data systems require many of them and, as a result, are prone to stragglers [16, 17]. They need to decompose stateful and stateless components with greater care, due to the inefficiencies of remote storage access [18, 19]. The unit prices are substantially higher, because providers provision serverless resources on behalf of the users to provide elasticity [20, 21].
Previous work shows the viability of systems built on serverless infrastructure for infrequent and interactive analytical workloads [22, 23]. Ana-logously, this workload category is also targeted by serverless databases that integrate automatic and elastic scalability at the system level [24, 25, 26].
In this paper, we present Skyrise, the first fully serverless query processor. Skyrise has no server-based component and scales down to zero when there is no load. Skyrise offers techniques to improve performance robustness and cost efficiency. We make the following contributions:
  1. (1)
    We present the Skyrise query processor, which is an open-source end-to-end SQL system that builds entirely on serverless infrastructure. [^1] Skyrise is available at https://github.com/hpides/skyrise.
  2. (2)
    We propose adaptive techniques for robust query performance on performance-variable cloud resources. Specifically, we describe the pre-warming of serverless resources and the re-triggering of straggling serverless query workers.
  3. (3)
    We introduce a cache for intermediate results backed by serverless storage, making Skyrise more cost-efficient for high query volumes.
  4. (4)
    We evaluate the performance, cost efficiency, and elasticity of Skyrise. We show that runtime and cost for TPC-H queries are on par with that of other cloud data systems.
The rest of the paper is structured as follows. In Section 2, we review background on serverless cloud infrastructure. Then, we give an overview of the architecture of Skyrise and explain key design decisions in Section 3. We then evaluate the performance, cost efficiency, and elasticity of Skyrise in Section 4. We discuss related work in Section 5 and conclude in Section 6.
2 Serverless Infrastructure
In recent years, serverless cloud infrastructure has matured into a set of reliable services which are effectively utilized in production settings. All major cloud providers have serverless compute (e. g., Azure Functions [15] and AWS Lambda [14]) and storage offerings (e. g., Amazon S3 [27] and Google Cloud Storage [28]).
The services provide elastic and automatic scalability through large multi-tenant pools of resources. Besides elasticity and resource management covered by cloud providers, billing of resources at fine consumption granularities allows for a pay-per-use (PPU) economic model. For that, providers need to match the demand closely and instantly to avoid provisioning/cost of idle capacity, and consequently provision resources on behalf of the users, which result in higher unit prices. To protect the service, quotas are enforced per user through admission control.
Subsequently, we discuss the two types of infrastructure that resemble core building blocks for data processing systems, namely function as a service (FaaS) platforms for computation and serverless storage services.
2.1 Function as a Service Platforms
Serverless function platforms like AWS Lambda are commonly realized by combining client-facing load balancing, cluster management, and functions / workers for which the system communicates through eventing. The cloud function user submits application binaries to be executed as ZIP archives or container images through a load balancing frontend service, coordinating function invocations. The request frontend checks the function metadata with the admission (control) service of the cluster management and—if the request does not exceed the user’s quota of concurrent function executions—it issues sandboxed environments from a pool of workers for the execution of the binary. Sandboxes are virtualized resources that are reused to place incoming requests, making them small, short, and ephemeral tasks.
Current vendors of FaaS platforms limit their configurations as shown in Table 1 with function sizing based on memory capacity that determines the number of virtual CPUs. Notably, function sizes are an order of magnitude smaller than VMs such as EC2 instances. For serverless functions hour-long lifetimes, persistent state, and direct communication are prohibited, making the functions small, short, ephemeral tasks that are easy to manage.
Table 1: Sizing and pricing of AWS compute services.
Resource Lambda (ARM) EC2 (C6g)
Memory Configurable Configurable
Capacity [GiB] 0.125 – 10 2 – 128
Price [¢/GiB-h] 3.84 – 4.80 0.65 – 1.70
Compute Memory-based Configurable
Capacity [vCPU] 0.07 – 5.79 1 – 64
Price [¢/vCPU-h] 6.79 – 8.49 1.30 – 3.40
Network Constant Compute-based
Bandwidth [Gbps] 0.63 0.375 – 25
Price [¢/Gbps-h] 0.48 – 0.60 3.27 – 8.70
For elastic scaling, serverless compute requires quick startup times of hundreds or even thousands of serverless functions. In Table 2, we depict the different startup times of a serverless function that was either freshly created and invoked (cold start) or created, invoked, and invoked again (warm). Cold starts show 20 to 50 times higher initialization latencies than warm starts, also with higher variation, which can be explained by the clean-up strategy of unused underlying virtual machines.
2.2 Serverless Storage Services
Cloud storage suitable for serverless compute requires elastically scalable capacity with high availability and durability to serve concurrent requests from a high number of serverless workers. The canonical choices are object stores (e. g., Amazon S3 [27], Azure Blob Storage [29]), key-value stores (e. g., DynamoDB [30], GCP Firestore [31]) or shared filesystems (e. g., EFS [32], Azure Files [33]). Object storage is designed to store immutable binary objects of varying sizes and formats at cloud network bandwidth. In contrast, Key-value stores and networked NFS/SMB file systems have lower latency key lookups at higher IOPS for kilobyte-sized values, but lower bandwidth and throughput.
Current storage infrastructures utilize similar architecture components for load balancing and rate limiting as serverless compute. The interaction is done via HTTPS requests that are evenly distributed across storage servers that hold partitioned data as shards. The economic model of cloud storage features static storage costs and dynamic costs per usage for requests and transfers, thus fitting well to the serverless PPU model.
Table 2: Startup latency of AWS compute instances.
Latency [ms] Min Max Avg
Lambda Cold Start 122 451 185
Lambda Warm Start 5 9 6
EC2 Cold Start 12,795 22,817 15,226
EC2 Warm Start 9,810 19,288 11,512
Disaggregated cloud storage exhibits a latency orders of magnitude higher than local storage, as shown in Table 3 for median and tail latencies. For one million 1 KiB read and write requests, S3 Standard shows the highest latencies of $27$ms read and $40$ms write latency. S3 Express trades availability for performance and achieves latencies comparable to DynamoDB.
Table 3: Performance and cost characteristics of AWS serverless storage services.
Service Median Latency Tail Latency Requests Transfers Storage
[ms] [ms] [¢/M] [¢/GiB] [¢/GiB-mo]
Read Write Read Write Read Write Read Write
S3 Standard 27 40 $>$1k 500 40 500 0 0 2.1–2.3
S3 Express 5 8 120 150 20 250 0.15 0.8 16
DynamoDB 4 6 100 250 25 125 0 0 25
EFS 6 15 100 600 0 0 3 6 16–30
From the cost perspective, S3 Standard is the most economical storage choice with the lowest static storage costs, no transfer costs, but the highest costs for number of requests. In contrast, S3 Express halves the request costs, but introduces transfer costs and seven times higher storage costs. The other storage options have factor around 10 higher storage costs compared to S3 Standard and either demand costs per request for DynamoDB or data transfers in case of EFS.
3 System Architecture
In this section, we introduce the Skyrise serverless query processor. We provide an overview of its architecture and describe the core components, which fall into three layers: Query compilation, query execution, and storage management.
Skyrise employs a distributed shared-storage architecture based on serverless object storage. Data referenced by queries always originate and materialize on object storage. The Skyrise query coordinator and worker nodes run as serverless functions. They run only when actively involved in query processing. This way, Skyrise fully exploits the elasticity of modern cloud infrastructure.
[…]
3.1 General Overview
Figure 1: Architecture of the Skyrise query processor showing the component interactions for processing a query.
The main components of Skyrise and their interactions during query processing are depicted in Figure 1. To process a query, a user sends a SQL query string wrapped inside a JSON request to an HTTP endpoint [37]. This triggers a serverless function running an instance of the Skyrise query coordinator. The coordinator gets the query string and manages the lifecycle of just this single query. Additional calls to the same endpoint do not lead to queries being queued. Instead, the queries are managed by separate coordinator instances and run concurrently. The coordinator then parses the SQL string and compiles an executable query plan (cf. Section 3.2). This plan contains pipelines of physical operators and dependencies between the pipelines. For each pipeline, the plan defines the number of fragments for parallel execution. Then, the coordinator schedules the pipelines for staged execution based on their dependencies (see Section 3.3). When a pipeline starts, a serverless query worker function is invoked for each fragment. The worker instances begin local execution by reading their data partitions from shared storage (Section 3.4). Then, they process their query fragments and write the results back on shared storage. Once all workers finish, the pipeline ends and subsequent pipelines are scheduled until the query completes.
3.2 Query Compilation
Figure 2: Query compilation pipeline showing the intermediate and final compilation artifacts and their consumption.
The query compilation pipeline in Skyrise takes a SQL query and generates a physical plan for distributed data-parallel execution by serverless query workers. The pipeline consists of multiple steps and is depicted in Figure 2. The first set of steps resemble the SQL frontend (in green). This component parses a query string and builds an abstract syntax tree (AST) that is further transformed into a logical plan for this query (LQP). Thereby, semantic types of columns in referenced tables are validated against an external database catalog. Then, a rule-based optimizer framework applies a range of conventional heuristics-based rules to the plan structure. These rules include predicate pushdown, subquery flattening, and join ordering. They only need simple statistics and are oblivious of Skyrise’s distributed execution environment. Next, the query optimizer (in blue) employs another set of rules to produce a physical query plan (PQP). The rules map logical operators to physical operators, e.g., a repartition join vs. a broadcast join. They identify query pipeline breakers in the plan to introduce shuffle points as necessary. And finally, they prepare the plan for serverless execution. This involves deciding on the number and size of workers based on the input data partitions, as well as picking a strategy for shuffling, e.g., staging to reduce cost [22, 23] or tiering to reduce runtime [38]. The final artifacts of query compilation are the per-pipeline PQPs parameterized with the input and output data partitions for the individual workers. These plan fragments are then scheduled for execution.
3.3 Query Execution
The execution of a query in Skyrise is orchestrated by the query coordinator (cf. Figure 1). Once a query is compiled into a set of interdependent pipeline PQPs, the coordinator schedules these PQPs stage-wise based on their dependencies. For each fragment of a pipeline, the coordinator invokes a serverless function running an instance of the Skyrise query worker. When the number of fragments is large and function invocation can dominate query runtime, the coordinator employs a two-level invocation procedure [22]. For $W$ fragments, it invokes $\sqrt{W}$ workers, each with a list of $\sqrt{W}$ fragments that in turn invoke $\sqrt{W}-1$ workers before executing their own fragment. The workers and the coordinator both invoke cloud functions asynchronously, decoupling their lifetime from the lifetime of their direct child functions. A function request is a query fragment serialized in JSON. Upon startup, a query worker first deserializes its invocation payload into the PQP form. Then, it schedules and executes the operators accordingly. The worker fetches its input data partitions from shared storage (cf. Section 3.4) and transforms them into an in-memory columnar format. From there, the execution follows a push-based, vectorized model. The query operators push their results in batches to downstream operators. Operators with no interdependencies run in parallel. Operators further exploit the thread-level parallelism available in serverless function workers. The final operator in a fragment materializes the results to shared storage. Before the worker quits, it sends a response message to a queue that is polled by the coordinator. The response contains the result location and execution statistics that can be used for adaptive execution behavior and query billing. When the coordinator has received responses from all workers in a stage, it proceeds with scheduling the next stage until the query finishes. Upon query completion, the coordinator responds to the user with the location of the final query result.
3.4 Storage Management
Figure 3: Storage I/O stack of the Skyrise query worker including the interactions between query operators and cloud storage.
Skyrise disaggregates storage from compute, and as a result, it is subject to the inefficiencies around remote storage access [18, 39]. In this section, we describe the Skyrise storage stack, which employs a number of techniques to mitigate these issues.
Skyrise uses serverless object storage [27, 28] for reading and writing table data during query execution. Thus, its storage stack is designed to efficiently operate on data stored in large objects and formatted in PAX-style [40] file formats, such as Parquet [19] and ORC [41]. The key techniques in this setting are right-sizing and parallelizing storage requests, as well as deeply integrating with the query execution and data format components.
Figure 3 shows the components in the Skyrise storage layer and how they interact with the query execution layer and external cloud storage. There are two dedicated handlers for storage I/O operations that are decoupled from the core query execution and backed by a separate thread pool. This separation accounts for the different resource demands of the two layers.
The input handler splits large object storage requests from the execution layer into several smaller requests, which are processed in parallel. The byte offsets of the small requests are chosen based on the object’s PAX layout. This way only relevant columns and rows are fetched. The result of each request is stored in a dedicated buffer. Straggling requests are retriggered after a timeout based on their size [23]. Once buffered, the results are deserialized and decompressed batch-wise and on demand, keeping the memory footprint of the I/O stack small. The output handler takes data batches from the execution layer and serializes, compresses, buffers them as they arrive. Once the query worker’s result is complete, it is materialized to remote storage as a single object.
4 Experimental Evaluation
In this section, we conduct a series of experiments to answer the following questions about the FaaS-based serverless query processor Skyrise:
  • How does Skyrise’s query performance compare to other cloud-based data processing systems?
  • Is Skyrise cost-efficient compared to commercial Query-as-a-Service (QaaS) systems?
  • How well does Skyrise scale elastically?
4.1 Setup and Methodology
We conduct all experiments on AWS in the region us-east-1 [42] and in the time frame of August to October 2024. For our experiments on Skyrise, we deploy ARM-based Lambda functions. Our driver has low resource requirements and runs continuously on an c6g.xlarge instance across experiments. All other compute resources are newly created for each experiment configuration and repetition. As cloud data processing systems, we consider AWS Athena [24], Snowflake [43], and Lambada [22]. We use queries and datasets from the TPC-H [44] benchmark. The tables are partitioned into 180 MB Parquet files using ZSTD compression and are stored on S3. We employ the standard generators and do not partition or sort on any specific keys. The lineitem table is 177.4 GiB in total and the orders table 44.9 GiB.
4.2 End-to-End Workloads
To assess the performance and cost efficiency of Skyrise, we compare it against several state-of-the-art approaches: The academic research prototype Lambada [22] and the commercial offerings Snowflake and Athena. Lambada (see Section 5) is a research prototype for FaaS-based query processing. We report the numbers published in [22]. Snowflake is a commercial cloud-based data warehouse offering. Typically, users allocate several warehouse instances. Recently, Snowflake added serverless tasks which are mainly aimed at data maintenance. Serverless tasks are executed on warehouses provisioned by Snowflake and are billed per second. Serverless tasks are created with an initial warehouse size provided by the user. After a few executions, Snowflake automatically determines the warehouse size for each tasks. For stable measurements, we created new tasks per query execution and set the initial size to 2xlarge (the configuration with the best price-performance ratio). We loaded the TPC-H dataset as external tables. We use Snowflake version 8.38.3 and take costs provided by Snowflake’s monitoring tools (with a standard configuration costing USD 2 per compute credit). Athena is a serverless query processing offering by Amazon. Similar to the Snowflake setup, we created external tables based on the TPC-H dataset. For query execution, Athena uses Trino (previously Presto). Resources used for queries are automatically determined by Athena and cannot be adapted or observed by users. The determine costs per query, we use the reported bytes scanned by Athena and the current pricing model of USD 5 per terabyte of data scanned.
We evaluate the three scan-heavy TPC-H queries 1, 6, and 12 (scale factor 1000), which are well-suited for serverless execution as they are not dominated by shuffling operators. We measure runtimes and costs per query execution, reporting the median of ten executions.
4.2.1 Query Latency
Figure 4: Query performance for TPC-H queries.
Users with ad-hoc analytical query workloads want interactive query performance. We present our results in Figure 4. We see a similar pattern for all queries. The fastest system is in all cases Lambada, followed by Skyrise, which is on average 30 % slower. The two commercial systems Athena and Snowflake are slower for evaluated queries with Athena being 5×5$\times$5 × faster than Snowflake in average.
We assume the reason for the performance difference between Lambada and Skyrise to be the dataset that was used for the Lambada experiments. This dataset differs in two significant ways. First, the string columns in the dataset were replaced with integers. Second, the lineitem relation is sorted on the l_shipdate attribute, which enables partition pruning for the employed queries.
We observe in separate experiments that both commercial systems perform better on smaller scale factors in their serverless modes, but their performance degrades on terabyte-scale queries.
4.2.2 Monetary Cost
Figure 5: Query costs for TPC-H queries.
Figure 5 shows the average costs per query. The pattern is comparable to the latency results. Lambada has the lowest costs per query, followed by Skyrise. Athena and Snowflake follow with Athena being 7×7$\times$7 × cheaper in average.
To conclude, Skyrise achieves comparable latencies and costs as the state-of-the-art system Lambada. The commercial system Athena closely follows the research systems, which is noteworthy given that the non-commercial systems Lambada and Skyrise only incur infrastructure costs. In our tested scenario, the recent Snowflake serverless tasks offering cannot compete with Athena.
4.3 Elasticity
Figure 6: Aggregated query latencies for TPC-H queries 1 & 6.
To evaluate the elasticity of Skyrise, we run TPC-H queries 1 and 6 for five different scale factors, ranging from 1 to 10 000. We run the queries cold with 15 minute intervals between runs to avoid any caching or provisioning effects in the AWS infrastructure. The results are shown in Figure 6. We report the aggregated runtimes of both queries. We observe that larger scale factors, especially 1 000 and 10 000 are generally more prone to stragglers in function invocations and storage requests. This is particularly the case for the 10 000 configuration, which runs with a maxmimum of 2,500 query workers and suffers from IOPS limits in S3. Generally, without any upfront provisioning or query load, Skyrise automatically scales the resources with the size of the query.
5 Related Work
In recent years, both academia and industry have been interested in serverless data processing, with a particular focus on elasticity. There are research prototypes that build upon FaaS platforms and serverless storage to inherit their elasticity. Other, primarily commercial, systems add techniques to reallocate resources between their tenants in cloud environments. We relate to both lines of research.
5.1 FaaS-based Data Processing
After its first public availability in 2014, FaaS technology was first prominently picked up for cloud computing by Jonas et al. who developed PyWren [16]. PyWren is a prototypical implementation of a programming model for disaggregated cloud computing environments using stateless functions.
While PyWren more generally targeted cloud applications, the potential of cloud functions for data analytics was shown by Kim et al. with Flint [45] and later Pu et al. with Locus [46]. Flint is a rewrite of Apache Spark’s execution layer using serverless building blocks like workers and disaggregated storage. In addition, Locus proposed a shuffling operator and intermediate data exchange using AWS ElasticCache and S3.
For data processing on serverless infrastructure, Locus was extended by Zhang et al. with a serverless SQL engine [47], called Caerus, focusing on analytical query processing atop AWS Lambda and cost-efficiency. The Lambada system [22] spawns static cluster for all query pipelines.
The shuffling approaches from PyWren, Flint, and [48] were further enhanced by Lambada and Starling [23] through stage shuffles, reducing shuffle I/Os via ephemeral storage. In addition Starling addresses robust performance processing data on S3 through pipelining and solutions for data skew and stragglers. Tiered shuffle was designed in Locus, Cackle [49]/Pixels [50], Caerus /Jiffy [51] and Pocket [52]. Boxer [53] shuffles via NAT and Flock [54] for streaming utilizes function invocation payloads [54].
Narasayya and Chaudhuri surveyed cloud data services and discussed the challenges for serverless computing [55]. The authors mention overheads to write worker state due to their short lifetimes or inefficiencies that arise due to missing network connections between workers. Workloads that require communication (e.g., transaction protocols or shuffles) thus need to write data to slow object stores. When adapting existing DBMS to serverless setups, the authors list assumptions in current systems that need to be reconsidered and the challenge to balance performance and costs of (almost) infinitely scalable processing capabilities.
Combining stateful and stateless compute is an interesting new area that Pixels [50] and Cackle [49] introduced as hybrid compute.
5.2 Elastic Cloud Data Processing
Presto [56] is an open source query engine for relational and non-relational data. Trino (previously known as PrestoSQL) is the backend of Amazon’s commercial serverless Athena [24] offering.
Redshift [11] is a distributed cloud-based data warehouse by Amazon. Its companion, Redshift Spectrum allows to run queries serverless directly on object store data.
Dremel is a distributed analytical query processing system developed by Google [12]. Dremel is the underlying technology behind BigQuery.
Snowflake [57] is a commercial cloud-based warehouse solution. Pricing is based on the user-selected warehouse size (per-second billing), storage usage, and egress data. Recently, Snowflake added capabilities to schedule serverless tasks (see Section 4).
Databricks introduced the concept of data lakehouses [58]. Lakehouses are cloud-based analytical systems similar to data warehouses like Snowflake with the main difference that they directly process queries on open direct-access data formats, such as Parquet. Users can directly query structured, semi-structured, and unstructured data without upfront data ingestion. Being initially built on Spark, the lakehouse architecture aims to combine machine learning with analytical query processing.
5.3 What is different in Skyrise?
In contrast to the related work, Skyrise is the first open-source, end-to-end SQL system built entirely on serverless infrastructure. The open challenge of robust query processing on flaky cloud resources is addressed by adaptive techniques (incl. pre-warming, retry for stragglers). For more cost-efficient processing of high query volumes, Skyrise introduces intermediate result caching. Having Skyrise open-source, allows for using it for research and teaching purposes at HPI and beyond.
6 Conclusion
In this paper, we present the serverless query processor Skyrise. Skyrise is completely built on the serverless premise and fully scalable from zero to thousands of cloud functions, with no stateful coordination. Skyrise is open-source and publicly available. Our experiments show Skyrise’s scalability and cost-efficiency for sporadic and variable workloads.
[…]
Supplementary information The source code, data, and/or other artifacts have been made available at https://github.com/hpides/skyrise.
[…]
Acknowledgements We thank the Master’s students that worked on the Skyrise project and implemented components of the query processor: Tobias Maltenberger, Julian Menzler, Theo Radig, Pascal Schulze, Jan Siebert, and many more.
This work was partially funded by SAP, the AWS Cloud Credit for Research Program, the German Research Foundation (414984028), and the European Union’s Horizon 2020 research and innovation programme (957407).